A box of inertia m is pulled with a constant acceleration a along a horizontal frictionless floor by a wire that makes an angle of 15° above the horizontal. If T is the tension in this wire, then what is T compared to ma? Is T=ma? Is T<ma? Is T>ma?
1
Expert's answer
2020-05-13T09:56:37-0400


Let's investigate forces, acting on a box. It's a gravity force, reaction force and tension force of the rope. Writing equation of motion in horizontal projection, we have: "T\cos{15^o}=m a". From this equation we obtain: "T=\dfrac{1}{\cos{15^o}}ma". As "\cos{15^o}<1", so "\dfrac{1}{\cos{15^o}}>1", thus "T>ma"